# Incentivizing Quality Text Generation via Statistical Contracts

### v1

@@ 5 Empirical Evaluation
We evaluate the empirical performance of our cost-robust contracts using LLM evaluation benchmarks. We compute binary and multi-outcome contracts for two distinct families of tasks based on evaluation scores from known benchmark datasets, optimizing the contract objectives set forth in Section 2.2. Our action space consists of the 7B, 13B, and 70B parameter model versions of the open-sourced Llama2 and CodeLlama LLMs [38, 31], which share the same architecture and hence similar inference costs. The benchmark data is used to create an empirical outcome distribution for each LLM in the action space. In both cases, contract optimization targets of the largest model variant (70B), which is the most performant and costly. Implementation details are provided in Section C.3, and code is available at: https://github.com/edensaig/llm-contracts.
@@
To estimate the inference costs of the language models, we leverage their open-source availability. We use energy consumption data from the popular Hugging Face LLM-Performance Leaderboard [22, 23], which we then convert to dollar values using conservative cost estimates. Estimation details are provided in Section C.1. As a first-order assumption of cost uncertainty, we assume that inference costs of alternative generators are bounded from below by the cost of the most energy-efficient alternative model ($c_{1}$), and bounded from above by the cost of the alternative model with the highest energy consumption ($c_{n-1}$).
@@ 5.1 Binary-outcome contracts across tasks (code generation)
Proof in Section B.8. Proposition 3 allows us to compare performance across different evaluation tasks without being sensitive to the choice of contract objective and constraints such as monotonicity.
@@ 5.2 Multi-outcome contracts
In practical applications, contracts with a simple functional form are often preferred since they are easier to comprehend. We compute optimal contracts with two types of simplicity constraints: monotone contracts (weakly increasing payout), and threshold contracts (full budget for all scores above a threshold, and zero otherwise). Results are presented in Figure 2 (Right), and Table 2. For the min-budget and min-variance criteria, monotone cost-robust contracts have an intuitive three-level structure: Zero pay for outputs with the lowest quality score, base pay for intermediate scores, and extra pay for outputs with the highest quality score. While threshold contracts may resemble current pricing schemes more closely (see Section C.1), monotone contracts enable a lower overall budget while still maintaining a simple functional form. For min-pay, monotone cost-robust contracts are in themselves threshold contracts, however they may deter risk-averse agents as they only pay for highest-quality outputs. In Section C.2, we additionally analyze non-monotone contracts, and show that further economic efficiency can be achieved by sacrificing simplicity.
@@
Table 2 compares cost-robust and cost-aware monotone contracts across different performance metrics. We observe that cost-robust contracts setting sacrifice a marginal increase in objective values: a $15\%$ increase in the min-pay objective, an $1.7\%$ increase if optimizing for budget, and $6.7\%$ increase when optimizing for minimum variance. We refer to Section C.2 for further analysis of cost-robustness in the non-monotone setting.
@@
Table 2: Monotone contracts curated on the MT-bench dataset. Costs are in units of \$/1M tokens, and are order-of-magnitude estimates based on typical prices of electricity - see section C.1. The numbers in bold denote the relative increase from the optimal monotone contract that the cost-robustness sacrifices in each setting. The percentages denote the price of cost-robustness: how much the objective values increase relative to the cost-aware setting.
@@ Acknowledgements.
The authors would like to thank Nir Rosenfeld, Stephen Bates, and Michael Toker for their insightful remarks and valuable suggestions. Eden Saig is supported by the Israel Council for Higher Education PBC scholarship for Ph.D. students in data science. Funded by the European Union (ERC, ALGOCONTRACT, 101077862, PI: Inbal Talgam-Cohen).
@@ Appendix A Additional Related Work
Appendix B Deferred Proofs
B.1 Linear programs and equivalent forms
@@
In this appendix we include linear programs (LPs) for optimizing contracts and hypothesis tests. Non-negativity constraints on the variables are ommitted where clear from context.
@@
$b$ $=(\underbrace{0,\dots,0}_{\text{$n-1$ times}},\underbrace{1,\dots,1}_{\text{$m$ times}},)\in\mathbb{R}^{n-1+m}.$
@@
B.2 Main lemmas
@@
These are vectors with uniform costs (up to $c_{n})$. Note that the costs in $c^{(1)}$ are (weakly) lower than those in $c$, and vice versa for $c^{(n-1)}$. Intuitively, since the agent gravitates towards lower costs, it is harder to incentivize the target action against lower costs. We formalize this as follows:
@@
Consider a distribution matrix $F$, and two cost vectors $c\leq\overline{c}\in\mathbb{R}_{\geq 0}^{n}$ satisfying $c_{n}={\overline{c}}_{n}$ (i.e., $c$ is dominated by $\overline{c}$). Then the sets of feasible solutions for contract design settings $(F,c)$ and $(F,\overline{c})$ satisfy ${\mathcal{T}}_{(F,c)}\subseteq{\mathcal{T}}_{(F,\overline{c})}$.
@@
Proofs appear in Appendix B.3, establishing also the other direction:
@@
B.3 Proofs of main lemmas
@@
The costs $c_{k}$ are assumed to be increasing in $k$, and therefore ${\overline{c}}_{i}-{\overline{c}}_{n}={\overline{c}}_{i}-c_{n}<0$ for all $i$. Moreover, as $c_{i}\leq{\overline{c}}_{i}$ for all $i<n$, the RHS of eq. 9 satisfies:
@@
B.4 Proof of main theorem
@@
B.5 Proof of properties of optimal cost-robust contracts
@@ B.6 MLR
B.7 Min-variance contracts
Claim 1.
$\min_{t\in\mathbb{R}_{\geq 0}^{m}}$ $t^{T}Vt$ (17)
Proof.
The variance of $t$ is given by:
$=\sum_{j}p_{j}\left(t_{j}-\sum_{k}p_{k}t_{k}\right)^{2}$
$\mathrm{Var}\left({t}\right)$ $=\left\lVert{Rt}\right\rVert$
$=t^{T}R^{T}Rt$
$=t^{T}Vt$
B.8 Two-outcome settings
Claim 2.
@@ B.8 Two-outcome settings
For all $i\in[n]$, let $F_{i}=\mathrm{Bernoulli}(p_{i})$. As $m=2$, any contract $t$ is a two-dimensional vector. By 2, 3, it holds that the optimal contract is of the form $t^{*}=(0,t_{1}^{*})$ for any of the three objectives. To prove that the optimal payment $t^{*}_{1}$ is the same for all objectives, observe that all objective functions are monotonically-increasing in $t_{1}^{*}$:
@@
Appendix C Experiments/Empirical Evaluation
C.1 Inference Costs
To calculate the costs for each model, we use energy data from the Hugging Face LLM Performance Leaderboard [^7] https://huggingface.co/spaces/optimum/llm-perf-leaderboard.. The energy efficiency for each model is expressed in the leaderboard in units of output tokens per kWH. To convert to actionable costs we assume a rate of .105 \$/kWH, aligning with conservative energy costs in the United States and giving us order-of-magnitude approximations of the actual inference costs. The inference costs are presented in units of \$/1M tokens. The leaderboard data was taken from the experiments on the A100 GPUs,and for each model, we took the GPTQ-4bit+exllama-v2 quantization benchmark. Table 3 shows the energy costs on the leaderboard for Llama2 and CodeLlama. We note that energy data was missing for CodeLlama-70B, so we extrapolated from Llama2-70B-chat.
@@ C.1 Inference Costs
Table 3: Energy costs for the Llama2 and CodeLlama model families.
@@
Table 4: Model Verbosities
@@
As alluded to in the paper introduction, current market pricing schemes for LLM generation involve pay-per-token rates for which the user pays regardless of the response quality. For open-sourced models such as Llama2, there exist API services to run model inference, such as AWS and Microsoft Azure.Alternatively, some pricing schemes behave as threshold contracts: an unsatisfied user may request from the API to regenerate a response free of charge, and hence will only pay if the response quality is above some threshold. For this reason threshold contracts can offer a "satisfaction guarantee" while retaining a simple form.
C.2 Multi-outcome contracts: Further Analysis

### v2

@@ 5 Empirical Evaluation
We evaluate the empirical performance of our cost-robust contracts using LLM evaluation benchmarks. We compute binary and multi-outcome contracts for two distinct families of tasks based on evaluation scores from known benchmark datasets, optimizing the contract objectives set forth in Section 2.2. Our action space consists of the 7B, 13B, and 70B parameter model versions of the open-source Llama2 and CodeLlama LLMs [38, 31], which share the same architecture and hence similar inference costs. The benchmark data is used to create an empirical outcome distribution for each LLM in the action space. In both cases, contract optimization targets of the largest model variant (70B), which is the most performant and costly. Implementation details are provided in Section E.3, and code is available at: https://github.com/edensaig/llm-contracts.
@@
To estimate the inference costs of the language models, we leverage their open-source availability. We use energy consumption data from the popular Hugging Face LLM-Performance Leaderboard [22, 23], which we then convert to dollar values using conservative cost estimates (see Section E.1). As a first-order assumption of cost uncertainty, we assume that inference costs of alternative generators are bounded from below by the cost of the most energy-efficient alternative model ($c_{1}$), and bounded from above by the cost of the alternative model with the highest energy consumption ($c_{n-1}$).
@@ 5.1 Binary-outcome contracts across tasks (code generation)
Proof in Section D.7. Proposition 3 allows us to compare performance across different evaluation tasks without being sensitive to the choice of contract objective and constraints such as monotonicity.
@@ 5.2 Multi-outcome contracts
In practical applications, contracts with a simple functional form are often preferred since they are easier to comprehend. We compute optimal contracts with two types of simplicity constraints: monotone contracts (weakly increasing payout), and threshold contracts (full budget for all scores above a threshold, and zero otherwise). Results are presented in Figure 2 (Right), and Table 2. For the min-budget and min-variance criteria, monotone cost-robust contracts have an intuitive three-level structure: Zero pay for outputs with the lowest quality score, base pay for intermediate scores, and extra pay for outputs with the highest quality score. While threshold contracts may resemble current pricing schemes more closely (see Section E.1), monotone contracts enable a lower overall budget while still maintaining a simple functional form. For min-pay, monotone cost-robust contracts are in themselves threshold contracts, however they may deter risk-averse agents as they only pay for highest-quality outputs. In Section E.2, we additionally analyze non-monotone contracts, and show that further economic efficiency can be achieved by sacrificing simplicity.
@@
Table 2 compares cost-robust and cost-aware monotone contracts across different performance metrics. We observe that cost-robust contracts setting sacrifice a marginal increase in objective values: a $15\%$ increase in the min-pay objective, an $1.7\%$ increase if optimizing for budget, and $6.7\%$ increase when optimizing for minimum variance. We refer to Section E.2 for further analysis of cost-robustness in the non-monotone setting.
@@
Table 2: Monotone contracts curated on the MT-bench dataset. Costs are in units of \$/1M tokens, and are order-of-magnitude estimates based on typical prices of electricity (see Section E.1). The numbers in bold denote the relative increase from the optimal monotone contract that the cost-robustness sacrifices in each setting. The percentages denote the price of cost-robustness: how much the objective values increase relative to the cost-aware setting.
@@ Acknowledgements.
The authors would like to thank Nir Rosenfeld, Ariel Procaccia, Stephen Bates, and Michael Toker for their insightful remarks and valuable suggestions. Eden Saig is supported by the Israel Council for Higher Education PBC scholarship for Ph.D. students in data science. This work received funding from the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation program (grant No.: 101077862, project: ALGOCONTRACT, PI: Inbal Talgam-Cohen), by the Israel Science Foundation (grant No.: 3331/24), by the NSF-BSF (grant No.: 2021680), and by a Google Research Scholar Award.
@@ Appendix A Additional Related Work
Appendix B Extensions
B.1 Targeting a set of high-quality models
For high-stake tasks such as summarizing medical information, it makes sense to target the most advanced model. In other scenarios, specifying a single target action can act as an intermediate step toward the final contract design.
For instance, a principal might aim to incentivize the use of any model that meets a certain quality threshold, e.g., requiring text generation from any LLM with more than 70B parameters. Formally, given a set of models ${\mathcal{G}}={\left\{{g_{1},\dots,g_{n}}\right\}}$ with associated costs $c_{1}<c_{2}<\dots<c_{n}$, we assume that higher-cost models offer higher quality. Let $k\in[n]$ represent the index of the minimum quality model that the principal seeks to target, such that the goal is to incentivize any model $g_{i}\in{\left\{{g_{k},\dots,g_{n}}\right\}}$.
To compute the optimal contract in this setting, the principal enumerates over the different target models $g_{i}\in{\left\{{g_{k},\dots,g_{n}}\right\}}$, designs an optimal single-target contract for each, and selects the “best” contract among the resulting designs – formally, the contract minimizing $\left\lVert{t}\right\rVert$ for the appropriate norm (see Section 2.2). Since all enumerated contracts are designed to satisfy incentive compatibility and potentially cost-robustness, the optimal contract among them also satisfies these properties.
Appendix C Contract Implementability
C.1 Conditions for implementablity
The implementatbility of min-pay contracts is discussed in [14, Appendix A.2], and the implementatbility of min-budget contracts is discussed in [32, Appendix B.3.1]. In both cases, the implementability of the contract design problems is characterized by the following condition:
Proposition 4 (Implementablity; [e.g., 14, Proposition 2]).
In a contract design problem $(F,c)$ with $n$ possible actions, an action $i\in[n]$ is implementable if and only if there is no convex combination of alternative actions that results in the same outcome distribution $\sum_{i^{\prime}\neq i}\lambda_{i^{\prime}}F_{i^{\prime}}=F_{i}$ but lower cost $\sum_{i^{\prime}\neq i}\lambda_{i^{\prime}}c_{i^{\prime}}<c_{i}$.
We note that Proposition 4 holds for all objectives $\left\lVert{t}\right\rVert$ described in eq. 1, as feasibility only depends on the incentive compatibility constraints. Adding to the result above, we show that implementability holds in the the intuitive case where the target model has the highest expected quality:
Proposition 5 (Highest expected quality implies implementability).
For a contract design problem $(F,c)$, denote the expected quality of action $i$ by $q_{i}=\mathbb{E}_{j\in F_{i}}{\left[{j}\right]}$. If $q_{i}<q_{n}$ for all $i<n$, then the contract is implementable.
Proof.
By contradiction. Assume that $q_{i}<q_{n}$ for all $i<n$ but the contract design problem is not implementable. By Proposition 4, there exists a convex combination of actions such that $\sum_{i<n}\lambda_{i}F_{i}=F_{n}$ and $\sum_{i}c_{i}<c_{n}$. Denote the convex combination of distributions by $F_{\lambda}$. By definition, it holds that:
$\mathbb{E}_{j\sim F_{\lambda}}{\left[{j}\right]}=\sum_{i<n}\lambda_{i}\mathbb{E}_{j\sim F_{i}}{\left[{j}\right]}=\sum_{i<n}\lambda_{i}q_{i}$
But from the infeasibility assumption $\sum_{i<n}\lambda_{i}F_{i}=F_{n}$, and therefore it also holds that:
$\mathbb{E}_{j\sim F_{\lambda}}{\left[{j}\right]}=\mathbb{E}_{j\sim F_{n}}{\left[{j}\right]}=q_{n}$
and therefore $\sum_{i<n}\lambda_{i}q_{i}<q_{n}$, which contradicts the initial assumption that $q_{i}<q_{n}$ for all $i$. ∎
C.2 Designing cost-robust contracts for strictly-intermediate target actions
Our main analysis technique for cost-robust contracts requires a separation between the interval covering the costs of target actions, and the interval covering the costs of alternative actions (see proofs of Lemma 2 and Lemma 3). This assumption does not hold when the principal targets a strictly intermediate model, formally $g_{i}\in{\mathcal{G}}$ such that $i<n$ and $c_{i}<c_{n}$. This captures scenarios where the principal seeks to incentivize text generation from a medium-sized model, but not from larger, costlier models that are available. In such instances, guaranteeing the implementability of cost-robust contracts requires additional assumptions.
To illustrate this, consider the following design setting, with $n=3$ actions and $m=2$ outcomes:
$\begin{array}[]{ll}F_{1}=(1,0)&c_{1}=0\\ F_{2}=(0,1)&c_{2}=1\\ F_{3}=(0.5,0.5)&c_{3}=10\end{array}$
Assume that the principal targets action $i^{*}=2$. In this setting, the cost difference is upper-bounded by $c_{i^{*}}-c_{i}\leq c_{2}-c_{1}=1$, and the cost-robust min-budget contract corresponding to the bound $b=1$ is given by Theorem 1:
$t=\left(0,\frac{b}{\left\lVert{{F_{3}}-{F_{2}}}\right\rVert_{\mathrm{TV}}}\right)=\left(0,\frac{1}{0.5}\right)=(0,2)$
This contract is incentive compatible with respect to the target action, yielding expected utilities $u_{A}(t,1)=0$ for action $1$, $u_{A}(t,2)=1$ for action $2$, and $u_{A}(t,3)=-9$ for action $3$.
However, consider the following setting, with identical target and different outcome distributions:
$\begin{array}[]{ll}F_{1}=(1,0)&c_{1}=0\\ F_{2}=(0.5,0.5)&c_{2}=1\\ F_{3}=(0,1)&c_{3}=10\end{array}$
In this setting, an incentive-compatible contract exists (e.g., $t=(0,4)$), and the same cost upper bound $b=1$ holds, however the cost-robust contract design problems are unfeasible by Proposition 4, as in the truncated contract design setting $\left(F=(F_{1},F_{2},F_{3}),c^{\mathrm{const}}=(0,1,0)\right)$ there exists a convex combination of alternative actions $\lambda_{1}=0.5,\lambda_{3}=0.5$ which satisfies $\sum_{i\in{\left\{{1,3}\right\}}}\lambda_{i}F_{i}=F_{2}$ and $\sum_{i\in{\left\{{1,3}\right\}}}\lambda_{i}c_{i}^{\mathrm{const}}=0<c_{2}=1$.
Appendix D Deferred Proofs
D.1 Convex programs and equivalent forms
@@
In this section, we include linear programs (LPs) for optimizing contracts and hypothesis tests. Non-negativity constraints on the variables are ommitted where clear from context.
@@
$b$ $=(\underbrace{0,\dots,0}_{\text{$n-1$ times}},\underbrace{1,\dots,1}_{\text{$m$ times}})\in\mathbb{R}^{n-1+m}.$
@@
Proposition 8.
$\min_{t\in\mathbb{R}_{\geq 0}^{m}}$ $t^{T}Vt$ (9)
Proof.
The variance of $t$ is given by:
$=\sum_{j}p_{j}\left(t_{j}-\sum_{k}p_{k}t_{k}\right)^{2}$
$\mathrm{Var}\left({t}\right)$ $=\left\lVert{Rt}\right\rVert$
$=t^{T}R^{T}Rt$
$=t^{T}Vt$
D.2 Main lemmas
@@
These are vectors with uniform costs (up to $c_{n})$. Note that the costs in $c^{(1)}$ are (weakly) lower than those in $c$, and vice versa for $c^{(n-1)}$:
$c^{(1)}\leq c\leq c^{(n-1)}$
Where the relation $\leq$ is defined element-wise. Intuitively, since the agent gravitates towards lower costs, it is harder to incentivize the target action against lower costs. We formalize this as follows:
@@
Consider a distribution matrix $F$, and two cost vectors $c\leq\overline{c}\in\mathbb{R}_{\geq 0}^{n}$ satisfying $c_{n}={\overline{c}}_{n}$ (i.e., $c$ is dominated by $\overline{c}$, and the cost of the target action coincides). Then the sets of feasible solutions for contract design settings $(F,c)$ and $(F,\overline{c})$ satisfy ${\mathcal{T}}_{(F,c)}\subseteq{\mathcal{T}}_{(F,\overline{c})}$.
@@
Proofs appear in Appendix D.3, establishing also the other direction:
@@
D.3 Proofs of main lemmas
@@
Let $\overline{c}$ be a cost vector satisfying $c\leq\overline{c}$ and $c_{n}=\overline{c}_{n}$. The costs $c_{k}$ are assumed to be increasing in $k$, and therefore ${\overline{c}}_{i}-{\overline{c}}_{n}={\overline{c}}_{i}-c_{n}<0$ for all $i$. Moreover, as $c_{i}\leq{\overline{c}}_{i}$ for all $i<n$, the RHS of eq. 10 satisfies:
@@
D.4 Proof of main theorem
@@
D.5 Proof of properties of optimal cost-robust contracts
@@
Remark 1 (Extending the proof of Theorem 2 to cost-robust min-pay contracts).
By Lemma 3, the min-pay contracts for the design settings $(F,c^{-a}),(F,c^{-b})$ defined in the proof depend linearly on $a,b$, respectively, and thus their expected pay is also linear in the bounds. The inclusion argument in eq. 14 holds for min-pay contracts as well, and thus an argument analogous to eq. 16 can be constructed for the ratio of expected payments. The rest of the proof follows similarly.
@@ D.6 MLR
D.7 Two-outcome settings
Proposition 9.
@@ D.7 Two-outcome settings
For all $i\in[n]$, let $F_{i}=\mathrm{Bernoulli}(p_{i})$. As $m=2$, any contract $t$ is a two-dimensional vector. By proposition 9, proposition 10, it holds that the optimal contract is of the form $t^{*}=(0,t_{1}^{*})$ for any of the three objectives. To prove that the optimal payment $t^{*}_{1}$ is the same for all objectives, observe that all objective functions are monotonically-increasing in $t_{1}^{*}$:
@@
D.8 Hardness of all-or-nothing cost-robust contracts
While cost-robust contracts can be computed in polynomial time by solving the corresponding convex programs (see Section D.1), restricting the functional form of the solution to have only two levels of payment entails computational hardness:
Definition 6 (All-or-nothing contract; [32]).
A contract $t$ has an all-or-nothing functional form if there exists $B>0$ such that $t_{j}\in{\left\{{0,B}\right\}}$ for all $j\in[m]$.
In [32, Theorem 3], it is shown that computing a min-budget all-or-nothing contract is NP-hard. We show that the same reduction is applicable for min-budget cost-robust contracts:
Proposition 11 (Hardness).
Computing a min-budget all-or-nothing contract is NP-hard.
Proof.
By reduction from 3SAT. Given a 3-CNF formula, we show that there exists a cost-robust contract design problem such that the required budget of an all-or-nothing cost-robust min-budget contract is below a certain threshold if and only if the formula is satisfyiable.
First, given a 3-CNF formula, apply the polynomial-time reduction described in [32, Appendix B.5.3] to construct a min-budget contract design problem $(F,c)$. Denote the target action of the design problem by $n\in[n]$. By the construction described in [32, Equation (27)], it holds that $c_{n}=1$ and $c_{i}=0$ for all $i<n$. Thus, the contract design problem tightly satisfies the cost difference upper bound $c_{n}-c_{i}\leq 1$. Moreover, since all alternative actions have the same cost $c_{i}=0$, the cost vector $c$ it is also the worst-case cost vector in the cost uncertainty set induced by the bound $b$ (by Lemma 1). Therefore, $(F,c)$ is also a cost-robust contract design setting for the bound $c_{n}-c_{i}\leq 1$.
By the proof of [32, Theorem 3], there exists a threshold $B_{0}$ such that the required budget of an all-or-nothing min-budget contract in the design setting $(F,c)$ is below a certain threshold if and only if the 3-CNF formula is satisfiable. ∎
Appendix E Experiments/Empirical Evaluation
E.1 Inference Costs
To calculate the costs for each model, we use energy data from the Hugging Face LLM Performance Leaderboard [^5] https://huggingface.co/spaces/optimum/llm-perf-leaderboard.. The energy efficiency for each model is expressed in the leaderboard in units of output tokens per kWH. To convert to actionable costs we assume a rate of .105 \$/kWH, aligning with conservative energy costs in the United States and giving us order-of-magnitude approximations of the actual inference costs. The inference costs are presented in units of \$/1M tokens. The leaderboard data was taken from the experiments on the A100 GPUs. For each model, we took the GPTQ-4bit+exllama-v2 quantization benchmark. Table 3 shows the energy costs on the leaderboard for Llama2 and CodeLlama. We note that energy data was missing for CodeLlama-70B, so we extrapolated from Llama2-70B-chat.
@@ E.1 Inference Costs
Table 3: Estimated energy costs for the Llama2 and CodeLlama model families, according to the methodology described in Section E.1.
@@
Table 4: Model verbosities (average output length) of the models in consideration. Since the values are of the same order of magnitude, we assume for simplification that the choice of model does not significantly affect verbosity (see Section E.1).
@@
As described in Section 1, current market pricing schemes for LLM generation involve pay-per-token rates for which the user pays regardless of the response quality. For open-source models such as Llama2, there exist API services to run model inference, such as AWS and Microsoft Azure. In other scenarios, some pricing schemes behave as threshold contracts: an unsatisfied user may request from the API to regenerate a response free of charge, and hence will only pay if the response quality is above some threshold. For this reason threshold contracts can offer a “satisfaction guarantee” while retaining a simple form.
E.2 Multi-outcome contracts: Further Analysis